# NCPP: Nova CPU Performance Predictor on a Novel Dataset

### v1

@@ Abstract
CPU performance prediction, which involves forecasting the performance scores of a CPU based on its hardware characteristics during its operation, is a critical technology for computational system design and resource management in the big data era. However, this research field currently faces two significant challenges. First, collecting real-world data is challenging due to the wide variety of CPU products on the market and the highly specialized nature of relevant hardware characteristics. In the research process, this field lacks a standard dataset with unified hardware characteristics, wide data coverage, and comprehensive benchmarks. Second, existing methods based on hardware simulation models or machine learning exhibit notable shortcomings, such as lengthy simulation test cycles and low prediction accuracy. To bridge these gaps, we first collect, preprocess, and standardize historical data from the 4th Generation Intel® Xeon® Scalable Processors across multiple benchmark suites to create a new dataset, named PerfCastDB. Subsequently, we design a deep learning based model called Nova CPU Performance Predictor (NCPP) as the baseline for this new dataset. The NCPP network is designed based on group attention mechanism. It effectively quantifies the implicit relationships between hardware characteristics within and across groups and comprehensively models the impact of various hardware characteristics on CPU performance prediction. We conduct comparative experiments using the proposed PerfCastDB dataset. Compared to existing approaches, NCPP achieves superior evaluation results, demonstrating its effectiveness. Furthermore, we have open-sourced part of the dataset and the NCPP network code to facilitate subsequent research. The resources can be accessed at https://github.com/xiaoman-liu/NCPP.
@@
CPU Performance Prediction, New Dataset, Group Attention Mechanism
@@ 1 Introduction
As the brain of a computer system, the CPU handles the assignment and processing of tasks, and manages operational functions used by all types of computers. In recent years, the explosive surge in the demand for computational resources from both academia and industry has led to increasingly stringent requirements for high-performance CPU. According to our internal dataset (Figure.1), numerous CPU models are designed and manufactured in recent years. However, the architecture of CPU has become increasingly complex, presenting significant challenges to CPU performance evaluation.
@@ 3.1 Data Collection
SPECrate2017_int_base 11 measure the integer computation capabilities of a CPU
SPECrate2017_fp_base 14 measure the floating-point computation capabilities of a CPU
MLC Latency 9 measure the latency for CPU to access data from cache or memory
MLC Bandwidth 9 measure the bandwidth for CPU to access data from cache or memory
Stream 4 measure sustainable memory bandwidth and computation rate for simple vector kernels
HPCG 1 measure the computational performance of solving sparse matrix equations
@@ 3.3.1 Feature Division
By leveraging data from these benchmark suites, we can construct a comprehensive profile of Intel SPR performance, encompassing both computational and memory-related aspects. To better understand the hardware characteristics within the dataset and enhance the model learning efficiency, we categorize these characteristics into four main groups. We will briefly discuss these groups below. For detailed data, please refer to the appendix.
@@ 4.2 Intra-group Attention Module
The attention calculation process can be expressed by:
atten $=\text{softmax}\left(\frac{\left(W_{Q}Q\right)\left(W_{K}K\right)^{\mathbf{T}}}{\sqrt{d_{k}}}\right)\left(W_{V}V\right)$ (3)
where $W_{Q},W_{K},W_{V}$ are trainable weights in the network and $Q$, $K$, $V$ are equal to the input data, $d_{k}$ denotes the embedding dimension.
@@
$FFN$ $=\operatorname{ReLU}\left(W_{1}X+b_{1}\right)$ (4)
where $W_{1},b_{1}$ are the weights and biases of the fully connected layer in FFN. After this step, the calculation can be formed as Equation.5:
atten $=\text{self - attention}(X)$ (5)
$LN_{1}$ $=\text{LayerNorm}(X+\text{atten})$
@@
Inter-group Attention Module The concatenated outputs from the intra-group attention modules are further processed through an inter-group attention module. Similar to the intra-group module, this module uses multiple attention heads to perform self-attention on the combined feature groups, followed by "Add & Norm" operations and "Feed Forward" layers. The final outputs from these attention heads are concatenated to form a unified feature representation.
@@ 6 Visualization Results
Figure.8 provides a comprehensive visualization of the attention matrix and aggregated feature importance for both the "Other Group" and "Character group" in the model trained with "SPECrate2017 FP" data. The attention matrix for the "Character Group" (Figure.8) illustrates the attention weights among 6 features, revealing multiple strong interactions, which underscores the importance of capturing these interactions for accurate predictions. The aggregated feature importance for the "Other Group" (Figure.8) shows that "TDP" has a slightly higher importance compared to "Power_freq," indicating that "TDP" plays a more critical role in the prediction. For the "Character Group" (Figure.8), the importance of features is more evenly distributed, with "Preset," "Microcode," and "CPU_Stepping" exhibiting higher importance, while "CPU_Family" and "OS" have relatively lower importance. This distribution suggests that multiple features within this group contribute significantly to the model performance, and their interactions are essential for capturing the underlying patterns in the data.
Figure.9 provides a comprehensive visualization of the attention matrix and aggregated feature importance for the "Memory Group" and "CPU Group" within the "SPECrate2017 FP" model. The attention matrix for the "Memory Group" (Figure.9) illustrates significant interactions between the seven features, crucial for understanding their collective influence on the model’s instance-wise predictions. In contrast, the "CPU Group" (Figure.9) shows multiple strong interactions among twenty features, underscoring the importance of capturing these interactions for accurate predictions. The aggregated feature importance for the "Memory Group" (Figure.9) reveals that "DIMM_rank", "Density", and "DIMM_Num" are the most critical feature, followed by "DIMM_Total" and "DIMM_Freq," while "Organization" and "CL" have lower importance. For the "CPU Group" (Figure.9), "AVX3_TurboFreq" exhibits the highest importance, with other features like "TMUL_P1Freq" and "Core_per_Socket" also contributing significantly.

### v3

# Personal Homepage

@@ Abstract
The server central processing unit (CPU) market continues to exhibit robust demand due to the rising global need for computing power. Against this backdrop, CPU benchmark performance prediction is crucial for architecture designers. It offers profound insights for optimizing system designs and significantly reduces the time required for benchmark testing. However, the current research suffers from a lack of a unified, standard and a comprehensive dataset covering various CPU benchmark suites on real machines. Additionally, the traditional simulation-based methods suffer from slow simulation speeds. Furthermore, traditional machine learning approaches not only struggle to process complex features across various hardware configurations but also fall short in achieving sufficient accuracy.
To bridge these gaps, we firstly perform a streamlined data preprocessing and reorganize our in-house datasets gathered from a variety CPU models of 4th Generation Intel® Xeon® Scalable Processors on various benchmark suites. We then propose Nova CPU Performance Predictor (NCPP), a deep learning model with attention mechanisms, specifically designed to predict CPU performance across various benchmarks. Our model effectively captures key hardware configurations affecting performance in across various benchmarks. Moreover, we compare eight mainstream machine learning methods, demonstrating the significant advantages of our model in terms of accuracy and explainability over existing approaches. Finally, our results provide new perspectives and practical strategies for hardware designers. To foster further research and collaboration, we open-source the model https://github.com/xiaoman-liu/NCPP.
@@
CPU Performance Prediction, New Dataset, Deep Learning, Group Attention Mechanism
@@ 1 Introduction
As the brain of a computer system, the CPU handles the assignment and processing of tasks, and manages operational functions used by all types of computers. In recent years, the explosive surge in the demand for computational resources from both academia and industry has led to increasingly stringent requirements for high-performance CPU. Figure 1 presents the numerous CPUs that have been newly designed and manufactured in recent years. However, the architecture of CPU has become increasingly complex, presenting significant challenges to CPU performance evaluation.
@@ 3.1 Data Collection
SPECrate2017_int_base 11 it measures the integer computation capabilities of a CPU
SPECrate2017_fp_base 14 it measures the floating-point computation capabilities of a CPU
MLC Latency 9 it measures the latency for CPU to access data from cache or memory
MLC Bandwidth 9 it measures the bandwidth for CPU to access data from cache or memory
Stream 4 it measures sustainable memory bandwidth and computation rate for simple vector kernels
HPCG 1 it measures the computational performance of solving sparse matrix equations
@@ 3.3.1 Feature Division
By leveraging data from these benchmark suites, we can construct a comprehensive profile of Intel SPR performance, encompassing both computational and memory-related aspects. To better understand the hardware characteristics within the dataset and enhance the model learning efficiency, we categorize these characteristics into four main groups. We will briefly discuss these groups below. For detailed data format, please refer to our open-source code [34].
@@ 4.2 Intra-group Attention Module
We take the feature of a instance X as the attention input, the self attention mechanism can be expressed by:
$Q^{(l)}$ $=X^{(l)}W_{Q}^{(l)}$ (3)
$K^{(l)}$ $=X^{(l)}W_{K}^{(l)}$
$V^{(l)}$ $=X^{(l)}W_{V}^{(l)}$
[…]
where $W_{Q}^{(l)}$, $W_{K}^{(l)}$, and $W_{V}^{(l)}$ are trainable weight matrices for layer $l$, $d_{k}$ denotes the embedding dimension. Then outputs of the self attention heads are concatenated and linearly transformed:
[…]
where
$\text{head}_{i}^{(l)}=\text{Intra-Self-Attention}(Q_{i}^{(l)},K_{i}^{(l)},V_{i}^{(l)})$ (5)
and $W_{O}^{(l)}$ is the linear transformation matrix after concatenation for layer $l$.
@@
$FFN$ $=\operatorname{ReLU}\left(W_{1}X^{(l)}+b_{1}\right)$ (6)
where $W_{1},b_{1}$ are the weights and biases of the fully connected layer in FFN. Then the normalization layer is introduced as:
[…]
@@
Inter-group Attention Module The concatenated outputs from the intra-group attention modules are further processed through an inter-group attention module. Similar to the intra-group module, this module uses multiple attention heads to perform self-attention on the combined feature groups, followed by "Add & Norm" operations and "Feed Forward" layers in Figure.4. The final outputs from these attention heads are concatenated to form a unified feature representation.
[…]
where
$\text{head}_{i}^{(l)}=\text{Inter-Self-Attention}(Q_{i}^{(l)},K_{i}^{(l)},V_{i}^{(l)})$ (9)
and $W_{O}^{(l)}$ is the linear transformation matrix applied after concatenation for layer $l$. The structure of "Inter-Self-Attention" is similar to that described in Equation 3.
@@ 6 Visualization Results
Figure.8 provides a comprehensive visualization of the attention matrix and aggregated feature importance for both the "Other Group" and "Character group" in the model trained with "SPECrate2017 FP" data. The attention matrix for the "Character Group" in Figure.8 illustrates the attention weights among 6 features, revealing multiple strong interactions, which underscores the importance of capturing these interactions for accurate predictions. The aggregated feature importance for the "Other Group" in Figure.8 shows that "TDP" has a slightly higher importance compared to "Power_freq," indicating that "TDP" plays a more critical role in the prediction. For the "Character Group" in igure.8, the importance of features is more evenly distributed, with "Preset," "Microcode," and "CPU_Stepping" exhibiting higher importance, while "CPU_Family" and "OS" have relatively lower importance. This distribution suggests that multiple features within this group contribute significantly to the model performance, and their interactions are essential for capturing the underlying patterns in the data.
Figure.9 provides a comprehensive visualization of the attention matrix and aggregated feature importance for the "Memory Group" and "CPU Group" within the "SPECrate2017 FP" model. The attention matrix for the "Memory Group" in Figure.9 illustrates significant interactions between the seven features, crucial for understanding their collective influence on the model’s instance-wise predictions. In contrast, the "CPU Group" in Figure.9 shows multiple strong interactions among twenty features, underscoring the importance of capturing these interactions for accurate predictions. The aggregated feature importance for the "Memory Group" in Figure.9 reveals that "DIMM_rank", "Density", and "DIMM_Num" are the most critical feature, followed by "DIMM_Total" and "DIMM_Freq," while "Organization" and "CL" have lower importance. For the "CPU Group" in Figure.9, "AVX3_TurboFreq" exhibits the highest importance, with other features like "TMUL_P1Freq" and "Core_per_Socket" also contributing significantly.
@@ 7 Conclusion
Declarations
  • Funding:This research was supported in part by Intel Corporation. The company may cover the publication fee for this manuscript.
  • Conflict of interest/Competing interests: The authors declare that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper. Although the author is an employee of Intel Corporation, the development of the model and the preparation of this manuscript were conducted independently of the company’s direct business interests.
  • Ethics approval and consent to participate: Not applicable. The dataset used for this research is proprietary to Intel Corporation and does not involve any human subjects or sensitive personal data that would require ethical approval. All procedures followed were in accordance with corporate policies and legal requirements.
  • Consent for publication: Not applicable. This manuscript does not contain any individual person’s data
  • Data availability: The dataset utilized for developing the CPU performance prediction model is proprietary to Intel Corporation and cannot be made publicly available due to confidentiality agreements and the sensitive nature of the business. The use of the data strictly adhered to the company’s data protection policies.
  • Materials availability: Not applicable. Although the model training process involved the use of computational resources, no physical materials were utilized in this research..
  • Code availability: The code for the CPU performance prediction model developed in this study is publicly available. The code can be accessed on GitHub: [https://github.com/xiaoman-liu/NCPP].
  • The author was solely responsible for the procesing of the data, development of the model, conducting experiments, and the drafting and revising of this manuscript. Colleagues within Intel Corporation provided insights by sharing background information on features and offering their perspectives on the features. However, the author independently carried out the whole work.